OpenStudy (anonymous):

how do you get the 10th term of a geometric sequence: 18, -6, 2, 2/3

OpenStudy (anonymous):

The first term is 18 and the common ratio is -1/3, so the sequence can be described by a_n=18*(-1/3)^(n-1) where a_1=18. a_10=18*(-1/3)^9

OpenStudy (anonymous):

thanks soooo much... greatly appriciated!